Problem 13.42

Part A

An irrigation canal has a rectangular cross section. At one point where the canal is 18.4 m wide and the water is 3.90 m deep, the water flows at 2.30 cm/s . At a second point downstream, but on the same level, the canal is 16.9 m wide, but the water flows at 10.6 cm/s . How deep is the water at this point?

Express your answer in meters to three significant figures.

ANSWER:

  m  
Answer #1

from continuity equation

A1V1 = A2V2

HERE A = L*D

18.4*3.9*0.023 = 16.9*D*0.106

D = 0.921 m